You are here: Foswiki>PartnerObserving Web>TechNightLogs>2009June13UT (21 Jul 2009, NormCushing)Edit Attach

Parasitic LBC-Red PartnerObserving in Parallel to IRTC LFBG for 13 June 2009 UT

LBC Observer: J. Hill (Tucson)
IRTC Observers: D. Thompson, M. Pedani (mtn)
Telescope Operator: D. Gonzalez Huerta

Penguins defeat Redwings 2-1 in Game 7 of Stanley Cup Finals!!!!!!

Summary

The main conclusion is that the correct value of RedZ22Zero is in the direction of moving from 2.55 to 2.05, but the amount to move has yet to be determined.

Details (times in UT)

02:30 Shutter doors open, but problems with the telescope because of an apparent failure in the emergency stop circuit.

04:35 Jon R. has installed a bypass jumper in the E-stop box (for the 5-10 floor loop) to get things running again.

04:38 Start mirror ventilation with mirrors about 1 degC below outside ambient of 8.5 degC.

04:40 First IRTC preset.

04:45 LBC server program on CMU doesn't want to connect to TCS, But TCSSendWavefront on lbtmu06 works OK.

04:50 Reboot CMU, and that cured the "Connect LBT" issues whatever they were.

04:57 dofpia, /redonly, /first with RedZ22Zero=1.55. It certainly is an adventure running dofpia while those IRTC guys are offsetting the telescope.

05:17 rV coarse superfoc (w/o Blue side) 051823 - 052009. This significantly overshot the Z22 zeropoint as the donuts are on the high end of the sequence now. But we may still have had mirror thermal issues going on.

05:23 Changed RedZ22Zero from 1.55 to 2.05

05:24 dofpia, /redonly This added +137 nm Z22 in two iterations.

05:31 rV coarse superfoc (w/o Blue side) 053045 -

05:34 rV coarse superfoc (w/o Blue side) 053439 -

05:36 Changed RedZ22Zero from 2.05 to 2.55.

05:37 dofpia, /redonly This added +117 nm Z22 in two iterations.

05:41 rV coarse superfoc (w/o Blue side) 054156 -

05:45 rV coarse superfoc (w/o Blue side) 054555 -

05:47 Changed RedZ22Zero from 2.55 to 3.05.

05:49 dofpia, /redonly This added +110 nm Z22 in two iterations.

05:54 rV coarse superfoc (w/o Blue side) 055430 -

We are seeing frequent (1 in 12 ?) problems with DX primary failing to make tolerance.

05:57 rV coarse superfoc (w/o Blue side) 055811 -

06:01 Changed RedZ22Zero from 3.05 to 3.55.

06:02 dofpia, /redonly - Converged in a single iteration after changing Z22 by only 2 nm. ??? These extrafocal pupils clearly have the outer edge too bright. Why didn't it make any change in the measurement?

06:07 dofpia, /redonly - Still nicely converged.

06:09 Changed RedZ22Zero from 3.55 to 1.55.

06:10 dofpia, /redonly - Changed RedZ22Zero by -584 nm in 6 iterations. (-150,-150,-126,-72,-39,-47) and bright ring at the edge of the pupil is completely gone.

06:21 rV coarse superfoc (w/o Blue) 062131 -

06:24 rV coarse superfoc (w/o Blue) 062459 -

Created RB_rV_LongCoarse_superfoc.ob with 8 exposures instead of 5. Starting at -0.40 mm

06:28 dofpia, /redonly - Converged in two iterations.

06:34 rv longcoarse superfoc (w/o Blue) 063442 - Manual abort for editting error in OB.

06:37 rv longcoarse superfoc (w/o Blue) 063746 -

Blue channel telemetry has failed.

06:42 rv longcoarse superfoc (w/o Blue) 064300 - PB Progress is counting 9/8, 10/8, 11/8 etc.

06:46 Dave and Marco are starting a collimation run, so my data collection is going to be less efficient.

06:47 Change RedZ22Zero from 1.55 to 2.05.

06:48 dofpia, /redonly - Seems slow (maybe was waiting for telescope on-source) - increased Z22 by 128 in two iterations.

06:56 dofpia, /redonly - Second iteration duplicated 065738 on lbtmu06, but 070007 had already appeared on lbtmu41.

07:08 rv longcoarse superfoc (w/o Blue) 070828 -

07:14 rv longcoarse superfoc (w/o Blue) 071423 - telescope slewing during this sequence

07:19 Change RedZ22Zero from 2.05 to 2.55.

07:20 dofpia, /redonly

07:25 Started creating IRAF display scripts in /Repository jdisplay155.cl, jdisplay205.cl

07:33 Turn Off, and restart LBC server.

The LBC code is still in Red-Only mode.

Set RedZ22Zero to 2.05 for now. Blue has to go the opposite way, so BlueZ22Zero changed from 1.90 to 2.3.

07:46 Increase DX Ki from 0.005 to 0.01.

07:55 John goes home.

-- JohnHill - 13 Jun 2009
Topic revision: r3 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback